HYPERFINE-RESOLVED SATURATION SPECTROSCOPY OF METASTABLE N_{2} IN THE (1-0) BAND OF THE B̃^{3}Π_{g} − Ã^{3}Σ_{u}^{+} ELECTRONIC TRANSITION

\underline{D. FORTHOMME}, C. MCRAVEN, G. E. HALL, T. J. SEARS, Chemistry Department, Brookhaven National Laboratory, Bldg. 555A, P.O. Box 5000, Upton, NY 11973, USA.

A spectrometer for measuring sub-Doppler saturation spectra in a discharge flow cell has been built and tested on the (1-0) band of the B̃^{3}Π_{g} − Ã^{3}Σ_{u}^{+} electronic transition of ^{14}N_{2} near 11250 cm^{−1}. A cw Ti:sapphire laser is stabilized relative to a single frequency HeNe laser and offset scanned with a RF synthesizer. An amplitude-modulated bleach beam is counter-propagated with a phase-modulated probe beam through the discharge cell. Absorption and dispersion phase signals are demodulated in a probe beam receiver, then demodulated at the lower bleach modulation frequency and finally averaged with lock-in detection at the audio frequency of the AC-discharge. Isolated hyperfine lines are observed with a FWHM of about 7 MHz and the splittings can be recorded with a precision of a few hundred kHz. Comparisons will be made with more extensive prior work by Giesen et al.^{a}^{b}

aH. Geisen, D. Neuschäfer and Ch. Ottinger, Z. Phys. D. 4 (1987) 263-290

bH. Geisen, D. Neuschäfer and Ch. Ottinger, Z. Phys. D. 17 (1990) 137-144